# A Survey on Large Language Model Acceleration based on KV Cache Management

### v2

@@ 7.1 Text Dataset
QA-SG NarrativeQA(LongBench) [247] Story 200 18,409 W F1 EN
QA-SG Qasper [260] Paper 1585 5001 W F1 EN
QA-SG Qasper(LongBench) [247] Paper 200 3619 W F1 EN
QA-SG MultifieldQA-en [247]
Paper, Legal,
Gov, Google
200 4459 W F1 EN
QA-SG MultifieldQA-zh [261]
Paper, Legal,
Gov, Google
200 6701 W F1 ZH
QA-SG QuALITY [262] Story, magazine 381 Doc, 6737 Q 4203 W EM EN
QA-MT HotpotQA [261] Wiki 112,779 1138 W EM, F1 EN
QA-MT HotpotQA(LongBench) [247] Wiki 200 9151 W F1 EN
QA-MT 2WikiMultihopQA [263] Wiki 192,606 Q 639 W EM, F1 EN
QA-MT MuSiQue [264] Wiki 24,814 1827 W F1 EN
QA-MT DuReader [265] Baidu 200,000 Q, 1,000,000 Doc 396 W BLEU, Rouge-L ZH,EN
QA+RET NewsQA(M4LE) [245] News - 3679 W Acc EN
QA+RET C3(M4LE) [245] Textbook - 3797 W Acc ZH
7.1.1 Question Answering (QA) Task
Dataset for this task usually consist of question-answer pairs, and documents that contains the answer to the question. For a model to run such task, documents and questions are usually used as the model input, while the output can differ greatly. Some datasets’ answers are closed-ended, meaning that the model should only output its answer in designated form, typically multiple choice answers, while the open-ended answers take a more free form. According to the number of documents involved in a question-answer pair, we can categorize QA task datasets into single-doc QA(QA-SG) and multiple-doc QA(QA-MT). The detailed statistics of the datasets for question answering are provided in Table XIII.
TABLE XIV: Text Dataset-Summarization. In the Avg. Len: average length, Tok: tokens; W: words. In the Instances column, Doc: documents, Q: questions, Inst: instructions. Particularly, SPACE has the concept of ’Entity’, and R/Ent stands for reviews per entity. Sum stands for summary. In the Metric column, EM: Exact Match. PM: Partial Match. Acc: Accuracy. For MultiNews, Rouge-SU skip bigrams when having a distance larger than 4 words. Particularly, LooGLE utilizes GPT-4 for its QA and summarization task, using it for answer’s semantic judgement.
Task Name Source Instances Avg Len Metric Lang.
SUM CNN/Dailymail [266] News 300,000 766 W Rouge-1/2/L EN
SUM XSum [267] News 400,000 431 W Rouge-1/2/L EN
SUM QMSum [268] Meeting 232 Meets, 1808 Q 9070 W Rouge-1/2/L EN
SUM MultiNews [269] News 51,216 5866 W Rouge-1/2/SU EN
SUM-QB+
Reasoning+
QA
LooGLE [251]
Papers, Wiki,
Movie, TV
776 Doc, 6448 Q
19,367 W
24,005 Tok
BLEU, Rouge, METEOR,
BERT, GPT4,
EM, PM
EN,ZH
SUM GovReport [270] Gov 19,466
9409.4 W
Rouge-1/2/L EN
SUM VCSUM [271] Meeting 239 14,107 Tok F1, Gold Rouge-1 ZH
SUM SummScreenFD [272] TV 269,000 6613 Tok Rouge EN
SUM BigPatent [273] Patent 1,341,362 3573 W Rouge-1/2/L EN
SUM SPACE [274] Review
50 Entities,
1,140,000 Reviews,
100R/Ent,
1050 Sum
15,532 W Rouge-1/2/L EN
SUM SQuALITY  [275] Story 625
5200 W
Rouge-1/2/L,
METEOR,
BERT
@@ 50 Entities,
SUM+RET CNNNews(M4LE) [245] News - 3754 W Rouge-L EN
SUM+RET CEPSUM(M4LE) [245] E-Commerce - 4003 W Rouge-L ZH
SUM+RET LCSTS(M4LE) [245] News - 4102 W Rouge-L ZH
SUM+RET NCLS(M4LE) [245] NCLS [276] - 3470 W Rouge-L EN,ZH
SUM+RET WikiHow [245] Wiki - 3514 W Rouge-L EN
SUM+RET News2016 [245] News - 3785 W Rouge-L ZH
SUM Pubmed(M4LE) [245] Medical 1267 3678 W Rouge-L EN
SUM BookSum(M4LE) [245] Book - 2643 W Rouge-L EN
SUM CNewsum(M4LE) [245] News 690 1883 W Rouge-L ZH
SUM CLTS+(M4LE) [245] News - 3158 W Rouge-L ZH
SUM Arxiv(M4LE) [245] Paper 1550 3748 W Rouge-L EN
  • Qasper [260] consists of 5049 questions based on 1585 papers on NLP. Question is from NLP practitioners that only have read the abstract and title of a paper, then another set of practitioners answer these questions by reading through the whole paper. The supporting evidences is provided correspondingly. Each instance of the dataset consists of a question, an answer, corresponding paper and supporting evidence. Instances built by LongBench [247] doesn’t require evidence.
  • HotpotQA [261] is a typical for a multi-doc QA dataset. It’s built based on Wikipedia, and each instance consists of multiple documents, a question, an answer and supporting facts. Supporting facts is a set of paragraph indexes, annotated manually.
  • AltQA [253] is based on google’s NQ [258] dataset. The answer are all numerical. The original document is ”altered” so that each occurrences of the numerical answer is different from the original document, so as to avoid data contamination from pretraining. This dataset is also used in BAMBOO [246] benchmark.
  • PaperQA and MeetingQA from BAMBOO [246] benchmark are question answering tasks in the form of multiple-choice. Each instance of the two datasets consists of question , evidence, answer and corresponding content.
  • NarrativeQA [259] uses complex narratives that are self-contained as input documents. Both books and movie scripts are used. For question construction, annotators are only given a story summary, and are asked to write questions based on it. For each story(1572 stories in total), about 30 question-answer pairs are constructed from each summary-story pair. Notably, because of the consistency in story context, the task can be simplified to selecting a correct answer from all answers that relates to the story.
  • MultifieldQA [247] is an original dataset from Longbench. Its contents covers scientific papers, legal documents, government reports and google results. The dataset has both Chinese and English version, and each instance consists of context built on documents, and a question-answer pair.
  • 2WikiMultihopQA [263] is a multi-document QA dataset built on Wikipedia and Wikidata. WikiData is a Knowledge Graph database, from which the author was able to extract the (subject entity, property, object entity) triple that corresponds to a Wikipidia document. These triples are used as evidences in each QA pair, as a way for model to show its inference process. The dataset consists of 192,606 questions in total.
  • Musique [264] is also a multi-document dataset(or multi-hop dataset, as the paper refers to). Its data is extracted from existing single-hop QA datasets. These single-hop QAs are then composed into multi-hop QA pairs. In addition, Musique add some unanswerable QA pairs in order to further test model’s ability. There are 24,814 answerable questions in Musique, and each answerable question corresponds to an unanswerable question.
  • DuReader [265] is a multi-document QA dataset, whose data is based on Baidu search results. It consists of 200,000 questions, 1,000,000 documents and 420,000 answers. Each instance contains a question, multiple possible answers(also possible to be empty), and multiple documents.
  • TriviaQA [254] is a multi-document reading comprehension QA dataset. All QA pairs are from 14 trivia websites, written by trivia enthusiasts. For each QA pair, 6 supporting documents(evidence) are provided, collected from Bing search API as well as Wikipedia. The total number of QA pairs is 95,956, with a total of 662,659 supporting documents, the average length of each document is 2895 words.
  • TOEFL(L-Eval) [244] collect lectures from the TOEFL Practice Online as context . Each instance consists of a long input of lectures, multiple instructions(questions) and corresponding answers.
  • Coursera(L-Eval) [244] is a dataset built on Coursera website. Similar to TOFEL, Each instance consists of a long input of lectures, multiple instructions and corresponding answers.
  • SFiction(L-Eval) [244] is based on scientific fictions, in which context real-world principles don’t apply. The questions contained in the documents ask the model to answer it based on either contextual information or real-world knowledge, as a way to test model hallucination.
  • LongFQA(L-Eval) [244] is an open-ended QA dataset on finance based on earnings call transcripts.
  • CUAD(L-Eval)[244] is drawn from the CUAD [257] dataset, which use legal contract as its context.
  • QuALITY [262] is a multiple-choice single-document QA dataset. It uses science fictions, magazine articles and nonfiction articles as input documents. The question is written by those that have read the full document. Each instance contains a document, a multiple-choice questions and corresponding answers. Notably, part of the questions are unanswerable.
  • NewsQA [245] and DuoRC [245] are English QA datasets, constructed from news and movie plots, respectively.
  • C3 [245] is a multiple-choice QA dataset, based on second-language Chinese exams.
  • NQ [258] is a QA dataset based on Wikipedia pages. Each instance(or example, as referred to in original paper) consists of a question, corresponding wikipedia page, a long answer and a short answer.
TABLE XV: Text Reasoning/Classification Datasets. CLS: Classification. In the Avg. Len: average length, Tok: tokens; W: words. In the Instances column, Doc: documents, Inst: instructions. In the Metric column, EM: Exact Match. Acc: Accuracy.
Task Name Source Instances Avg Len Metric Lang.
CLS/Reasoning Long ListOps  [248] Generated 100,003 3106 W Acc EN
Reasoning ContractNLI  [277] Legal 10,319 2254 Tok EM EN
CLS LSHT(LongBench) [247] News 200 22,337 W Acc ZH
Reasoning GSM(16 shot) [244] GSM8K [278] 100 Doc, 100 Inst 5557 Tok Rouge-L, GPT-4, $\Delta L$ EN
Reasoning SenHallu(BAMBOO)  [246] Paper 200/200 3170/6357 Tok Precision, Recall, F1 EN
Reasoning AbsHallu(BAMBOO)  [246] Paper 200/200 3314/6445 Tok Precision, Recall, F1 EN
CLS MNDS News  [279] News 10,917 637 W Acc EN
7.1.2 Text Summarization Task
A summarization dataset is a curated collection of texts and their corresponding summaries. They typically include diverse content, such as news articles, scientific papers, or conversational data, paired with concise and accurate summaries. The detailed statistics of the datasets for text summarization are listed in Table XIV.
  • CNN/Dailymail [266], GovReport [270], and XSum [267] include a document and its corresponding summary in each instance. CNN/Dailymail is based on over 300,000 news articles, GovReport is based on 14,466 long government reports, and XSum is based on BBC news.
  • MultiNews [269] is a multi-doc summary dataset, each instance consists of multiple news and a summary.
  • Loogle [251] is based on papers, WikiPedia, movie and TV scripts. Each long input text corresponds to mutiple question-answer-summary triad. In total there are 776 documents and 6,448 questions. Average document length is 19.367 words.
  • VCSUM [271] is based on real-world Chinese meeting transcripts. Each meeting tarnscript corresponds to a headline, segmentation summaries and an overall summary. There’re 239 meetings in total.
  • SummScreenFD [272] is based on TV transcripts. Each instance consists of a TV transcript containing conversations, scenes and actor actions, and a summary(recapitulation, as referred to in original paper).
  • BigPatent [273] is based on 1,341,362 patent documents. The highlight of this dataset is that important information is distributed evenly in patent documents, compared to other types of documents. Each instance contains a document and its corresponding summary(human written abstract).
  • SPACE [274] is based on reviews of 50 hotels. The highlight of the dataset is that the summaries are written in 6 different aspects, based on the hotel’s review. Each hotel constructs an instance, containing the hotel’s name, multiple reviews, summaries of different aspects and an overall summary.
  • SQuality [275] is based on the same stories domain as QuALITY [262] dataset. It’s a query-based summarization dataset. Each instance contains a story, multiple summarization questions, and multiple summarizations that corresponds to each questions. There are 625 QA pairs in total.
  • CNNNews(M4LE) [245] is based on CNN English news. Each instance of the dataset is paired with a multi-sentence summary.
  • CEPSUM(M4LE) [245] is based on product information from Chinese e-commerce platform. Each instance contains a product description and corresponding summary.
  • LCSTS(M4LE) [245] is a summarization dataset in Chinese. It consists of over 2 million posts from a Chinese micro-blogging website, each post is paired with a summary. M4LE selects instances whose article has over 30 words.
  • NCLS(M4LE) [245] is a summarization dataset with articles and corresponding summaries in different language, which highlights model’s cross-lingual ability. Original NCLS is constructed from CNNNews and LCSTS.
  • WikiHow(M4LE) [245] is based on procedural descriptions on Wikipedia. Each article is entitled with a beginning of ”How to…”. Each paragraph of the article describes one step in the procedure, and corresponds to short summary. These summaries are then put together as the suymmary of the article.
  • News2016(M4LE) [245] is based on ove 2 million news articles in Chinese. For each article, its title is used as golden summary. M4LE remove instances whose length is less than 200 words or over 800 words.
  • PubMed(M4LE) [245] is based on medical papers. In M4LE, each paper’s abstract is used as the summary of the paper.
  • BookSum(M4LE) [245] is a dataset containing 405 English books, whose contents covers plays, novels and short stories. Each chapter of the content corresponds to a human-written summary.
  • CNewsum(M4LE) [245] is based on 304,307 news articles in Chinese. Each article corresponds to a human-written summary.
  • CLTS+(M4LE) [245] is based on CLTS [280]. CLTS contains over 180,000 Chinese articles, and CLTS+ uses back translation to make summaries more abstractive. M4LE selects part of these instances for benchmark.
  • Arxiv(M4LE) [245] is based on papers collected from arXiv.org. For each paper, its abstract is used as golden summary.
7.1.3 Text Reasoning Task
A reasoning task involves the ability of a model to draw logical conclusions, make inferences, or solve problems based on given information. It requires understanding relationships, patterns, or rules within the data to arrive at accurate and coherent outcomes.Natural Language Inference(NLI) can be considered a subset of reasoning. It highlights model’s ability to perform logical inference instructed by natural language.In an NLI task, the typical goal is to determine the relationship between two pieces of text: a premise and a hypothesis. The detailed statistics of the datasets for text reasoning are listed in Tab. XV.
  • Long Listops [248] is a mathematical reasoning dataset. It inputs an listop expression, instructing the model to perform calculation and output the exact numeric answer. A listop expression has a hierarchical structure that involves a set of simple mathematical operators. The final answer is a number in 0-9, described in original paper as ”a ten-way classification task”.
  • GSM [278] is a mathematcal reasoning dataset, which describes mathematical problems in natural language and ask the model to solve it.
  • ContractNLI [277] uses contracts as context, and provides hypothesis, answer, and added evidence to each instance as well. The task requires model to judge the relationship between the hypothesis and context. Each instance contains 607 contracts, each contract has 17 annotated hypothesis and corresponding answers.
  • LSHT(LongBench) [247] is a Chinese classification dataset. It’s based on Xinhua News. The model is asked to classify the input news articles into different categories.
  • SenHallu [246] and AbsHallu  [246]use content and a related hypothesis as model’s input, and instruct the model to determine whether the hypothesis is true based on the content. The false hypothesis(hallucination, as referred to by original paper) is generated by GPT.
  • MNDS News [279] is a classification dataset consisting of 10.917 news articles. The news articles have 17 first level categories and 109 second-level categories.
TABLE XVI: Text Dataset-Retrieval. In the Avg. Len: average length, W: words. Particularly, LongEval, StreamingEval and TopicRet is more of a data generation method, which makes their length and instance number flexible, denoted by ’-’. In the Metric column, Acc: Accuracy. F1 [281] calculates unigram overlap between model output and answers after processing elments like white-spaces and stop-words.
Task Name Source Instances Avg Len Metric Lang.
CLS/RET TREC(LongBench) [247] Web Question 200 5177 W Acc EN
RET LongEval  [252] Conversations - - Acc EN
RET StreamingEval  [136] LongChat[252] - - Acc EN
RET TopicRet(L-Eval)  [244] LongChat[252] - - Acc EN
RET DRCD(M4LE) [245] Wiki - 3617 W Acc ZH
CLS+RET MARC [245] E-Commerce 2200 3543 W F1 EN,ZH
CLS+RET Online Shopping(M4LE) [245] E-Commerce 2200 3714 W F1 ZH
CLS+RET MNDS News(M4LE) [245] MNDS News [279] - 3805 W Acc EN
CLS+RET THUCNews(M4LE) [245] News - 3721 W Acc ZH
7.1.4 Text Retrieval Task
A retrieval task in LLM benchmarks evaluates a model’s ability to retrieve relevant information from a large collection of data based on a given query. It tests the model’s understanding of the query, semantic matching, and efficiency in identifying the most relevant documents or pieces of information. The detailed statistics of the datasets for text retrieval are listed in Table XVI.
  • LongChat [252] has two subtask dataset for retrieval. Coarse-grained Topic Retrieval dataset use a long document that talk about a number of different topics, and instrutct the model to retrieve the first topic of the document. Fine-grained Line retrieval, on the other hand, is more challenging, which present the model with multiple lines that contain a diffrernt number and label, with similar line patterns. The model is asked to retrieve the number of a specific labeled line.Notably, such dataset can be easily constructed or generated, so it’s easy to create an ultra long dataset of this type. Because the dataset is easily constructed by definition, the length of the dataset and the number of instances is indefinite.
  • StreamingEval[136] construct a line retrieval task based on LongChat, which makes a query in every 10 lines, with its answer about 20 lines above, so as to evaluate the streaming conversation scenario.
  • TopicRet [244] on the other hand, is based on the coarse-grained topic retrieval task, but ask about the second or third topic instead of the first one, so as to make the task more challenging.
  • DRCD(M4LE) [245] is a reading comprehension dataset. In M4LE, DRCD is constructed into two subset, one(DRCD explicit) require model to return the articles’ IDs related to a given topic, and another subset(DRCD semantic) requires the model to answer specific questions given multiple paragraphs.
  • MARC [245] consists of bilingual(namely English and Chinese) reviews. The model is asked to identify all positive reviews and retrieve them.
  • Online Shopping(M4LE) [245] is based on 60K product reviews on Chinese e-commerce platforms. Reviews are categorized into positive and negative.
TABLE XVII: Text Dataset-Generation.In the Avg. Len: average length, Tok: tokens; W: words. In the Instances column, Doc: documents, Inst: instructions. In the Metric column, EM: Exact Match. Acc: Accuracy.
Task Name Source Instances Avg Len Metric Lang.
GEN LCC  [282] Code 360000 1337 W EM, Edit Sim Python/CSharp/Java
GEN RepoBench-P(LongBench) [247] Code 500 4206 W Edit Sim Python/Java
GEN/RET MultiDoc2Dial [283] Doc2Dial [284]
488 Doc,
4796 Dialogues
4283 T
F1, EM, SacreBLEU, Recall
@@ 7.1.4 Text Retrieval Task
GEN OpenReview(L-Eval)  [244] ASAP-Review[285] 20 Doc 60 Inst 11,170 Tok Rouge-L, GPT-4, $\Delta L$ EN
GEN ASAP-Review  [285] Paper
8877 Papers,
25,986 Reviews
6782 W/Paper
Rouge-1/2/L, BERT EN
GEN ShowsPred  [246] TV Shows 100/100 2389/4860 Tok Acc EN
GEN MeetingPred  [246] Meeting 100/100 3689/11578 Tok Acc EN
GEN-Code PrivateEval  [246] Code 152/152 3149/6230 Tok Pass@1 EN, Python
GEN-Code CodeU(L-Eval)  [244] Code 90 Doc 10 Inst 31,575 Tok Rouge-L, GPT-4, $\Delta L$ Python
TABLE XVIII: Text Dataset-Aggregation. In the Avg. Len: average length, Tok: tokens; W: words. In the Instances column, Doc: documents, Inst: instructions. In the Metric column, Acc: Accuracy. ES: Exponential Similarity, CI: Concordance Index
Task Name Source Instances Avg Len Metric Lang.
AGG SpaceDigest  [250] Reviews 500 5481 W ES EN
AGG BookSumSort  [250] Literature 500 6840 W CI EN
AGG PassageRetrieval-en [247] Wiki 200 9289 W Acc EN
AGG PassageRetrieval-zh [247] C4 Dataset 200 6745 W Acc ZH
AGG PassageCount  [247] Wiki 200 11,141 W Acc EN
AGG ShowsReport(BAMBOO)  [246] TV Shows 200/200 2992/6411 Tok CI EN
AGG ReportSumSort(BAMBOO)  [246] Reports 150/150 3753/8309 Tok CI EN
7.1.5 Text Generation Task
Generation tasks require model to generate contents based on the given instructions and context. The detailed statistics of the datasets for text generation are listed in Table XVII.
  • MultiDoc2Dial [283] gives model a dialogue history and all involved documents, and instruct model to generate the next turn of the dialogue.
  • OpenReview(L-Eval) [244], which is based on ASAP-Review [285], provides LLM with a paper and instruct it to generate a review.
  • ShowsPred and MeetingPred [246] use dialogue history as input, and ask model to infer which role said the last turn of the conversation. Apart from natural language context, code generation is also an important implementation for LLMs.
  • LCC [282] gives model long code snippets as context, and instruct model to generate the following line of code.
  • RepoBench-P [286] requires model to retrieve toe most relevant code snippets from a long input, and then generate code according to the instruction.
  • PrivateEval [246] use API documents and a code snippet as input, and instruct the model to generate code acccordingly. Notably, to avoid data contamination caused by pre-training, the keywords in API documents are modified, making the document ”private”.
  • CodeU [246] use the same practice of modifying keyword, only that it uses modified source code of public library, rather than API document, as an input.
7.1.6 Aggregation Task
Aggregation task involves understanding and aggregating information from the whole input to answer complex instructions, such as calculating the percentage of positive comments given a set of comments of different attitudes. The detailed statistics of the datasets for text aggregation are listed in Table XVIII.
  • SpaceDigest [250] give the model a set of hotel reviews, and ask the model to output the percentage of positive reviews in the context.
  • BookSumSort [250], ReportSumSort [246], and ShowsSort [246] use shuffled paragraphs from book summaries, TV transcripts or government reports as context, and ask the model to sort them in the correct order.
  • PassageCount [247] selects multiple passage, duplicates some of the paragraphs, and put all those paragraphs into an instance after shuffling. The model is then asked to determine how many documents are used to construct this instance.
  • PassageRetrieval [247], on the other hand, selects 30 wikipedia passages, and use GPT-3.5-Turbo to write a summary for one of them. Then these passages and the generated summary are used as the model input. The model is then instructed to tell which passage was the summary generated from.
7.1.7 Evaluation Metric for Text Datasets
General evaluation metrics used by text datasets mentioned above include Exact Match [287], Partial Match, Accuracy, Recall, Precision, F1, BLEU [288], SacreBLEU [289], Rouge [290], METEOR [291], BERT [292], Edit Similarity, Pass@k [293] , Exponential Similarity, Concordance Index, Mean Reciprocal Rank. In addition to general evaluation metrics, some more specific metrics are used in particular benchmarks. For datasets from L-Eval [244], the GPT-4 metric means the win-rate against Turbo-16K, judged by GPT-4. $\Delta L$ is the length difference between answer length and ground truth. For LooGLE [251], it utilizes GPT-4 for its QA and summarization task, using it for answer’s semantic judgment.
  • Exact Match (EM) [287] is a metric used to evaluate the accuracy of models in tasks like question answering or text generation. It measures the percentage of predictions that exactly match the ground truth answer, considering both the content and format.

### v3

@@ 7.1 Text Benchmarks
MMBench [279]
EN/ZH
MileBench [280]
@@
MLVU [281]
@@
LongVideoBench [282]
EN
Video-MME [283]
EN
NExT-QA [284]
EN
MVBench [285]
EN
MSVD-QA [286]
EN
MSRVTT-QA [286]
EN
  • MultiTurnBench [1] is a multi-turn long-context benchmark that includes 11 diverse datasets with varying query positions and multi-turn dependencies, solving the problem of unrealistic single-turn, end-of-sequence query evaluation by enabling realistic and challenging assessment of LLM performance in practical dialogue scenarios, which is built on existing benchmarks [273, 287, 288, 289, 290, 291, 292, 293, 294, 295].
  • NumericBench [267] is a benchmark designed to evaluate the fundamental numerical reasoning capabilities of LLMs, emphasizing their ability to understand the meaning of numbers.
  • RULER [268] is a synthetic benchmark designed to evaluate the long-context capabilities of LLMs across diverse task categories, including retrieval, multi-hop tracing, aggregation, and QA.
  • OneRuler [269] is a multilingual benchmark that evaluates long-context language models across 26 languages.
  • L-Eval [270] is a benchmark designed to evaluate long-context language models (LCLMs) across 20 diverse tasks, encompassing both closed-ended scenarios (e.g., reasoning) and open-ended ones (e.g., summarization). The benchmark includes datasets such as QuALITY [296], SPACE [297], among others.
  • M4LE [271] introduces a new benchmark for LLMs’ long-context comprehension, covering 36 datasets across 11 tasks and 12 domains, with input lengths from 1K to 8K words. It contains datasets such as TriviaQA [298], HotpotQA [287], BIGPATENT [299], MNDSNews [300], among others.
  • BAMBOO [272], evaluates LLMs’ long-text understanding across 5 tasks, including QA, hallucination detection, text sorting, language modeling, and code completion.
  • LongBench [273] is a comprehensive bilingual benchmark designed to evaluate the long-context understanding capabilities of large language models. It incorporates a diverse range of datasets, including HotpotQA [287], 2WikiMultihopQA [288], MuSiQue [289], Qasper [290], NarrativeQA [291], QMSum [292], GovReport [293], MultiNews [294], and TriviaQA [295], among others.
  • SCROLLS [274] is a benchmark designed to evaluate models’ ability to process naturally long texts across diverse domains such as literature, science, and entertainment. It includes tasks like summarization, question answering, and natural language inference, emphasizing the synthesis of information across extended inputs.
  • ZEROSCROLLS [275] serves as a benchmark aimed at assessing the zero-shot reasoning abilities of language models over extended texts.
  • LooGLE [276] benchmarks LLMs’ long-context reasoning using post-2022 documents, emphasizing reasoning over memorization, though future training may impact zero-shot evaluation.
  • LongEval [277] is a benchmark for assessing LLMs’ long-context capabilities, focusing on tasks requiring reasoning over extended text inputs.
  • StreamingEval [141] is a benchmark designed to assess the ability of instruction-tuned LLMs to perform question answering in streaming contexts, where input sequences grow continuously.
7.2 Multi-modal Benchmark
Multi-modal datasets combine various data types, such as text, images, audio, and video, to capture the complexity of the real world. Detailed discussions are provided for each benchmark, as outlined in Table XIV.
  • LLaVA-Bench [278] is structured around image-ground-truth textual description-question-answer triplets, segmented across COCO and In-The-Wild datasets.
  • MMBench [279] serves as a bilingual multi-modal benchmark, facilitating a comparative analysis of VLM performance across English and Chinese linguistic contexts.
  • MileBench [280] rigorously evaluates the multi-modal long-context capabilities of LLMs, including both diagnostic and realistic evaluation sets. It emphasizes long-context and multi-image tasks.
  • MLVU [281] is a comprehensive benchmark for evaluating multi-modal LLMs’ video comprehension, with longer videos and varied assessment tasks.
  • LongVideoBench [282] offers a framework aimed at assessing the capacity of large multi-modal models to comprehend lengthy videos with subtitles.
  • Video-MME [283] comprehensively evaluates large multi-modal models’ video analysis using 900 videos across diverse domains, ranging from 11 seconds to 1 hour for broad scenario coverage.
  • NExT-QA [284] notably boasts a dataset with 5,440 videos and approximately 52K manually annotated question-answer pairs, sorted into causal, temporal, and descriptive categories.
  • MVBench [285] comprises 200 multiple-choice question-answer (QA) pairs for each of the 20 temporal understanding tasks, amassing a total of 4,000 QA pairs.
  • MSVD-QA [286] is a collection of 1,970 video clips with descriptive captions, initially for video captioning.
  • MSRVTT-QA [286] comprises 10,000 video clips with 20 human-transcribed sentences each, focusing on connecting video content with language descriptions.
7.3 Evaluation Metric
In evaluating LLMs, multiple metrics are essential to comprehensively assess performance across various tasks and application scenarios. Below is a list of commonly used evaluation metrics with brief descriptions.
  • Exact Match (EM) [301] is a strict metric assessing model accuracy by requiring predictions to exactly match the ground truth.